### v1

# cNorthern Care Alliance NHS Foundation Trust (NCA), Manchester, UK

Abstract
: Accurate segmentation and classification of brain tumors from Magnetic Resonance Imaging (MRI) remain key challenges in medical image analysis, largely due to the lack of high-quality, balanced, and diverse datasets. In this work, we present a new curated MRI dataset designed specifically for brain tumor segmentation and classification tasks. The dataset comprises 6,000 contrast-enhanced T1-weighted MRI scans annotated by certified radiologists and physicians, spanning three major tumor types—glioma, meningioma, and pituitary—as well as non-tumorous cases. Each sample includes high-resolution labels and is categorized across axial, sagittal, and coronal imaging planes to facilitate robust model development and cross-view generalization. To demonstrate the utility of the dataset, we propose a transformer-based segmentation model and benchmark it against established baselines. Our method achieves the highest weighted mean Intersection-over-Union (IoU) of 82.3%, with improvements observed across all tumor categories. Importantly, this study serves primarily as an introduction to the dataset, establishing foundational benchmarks for future research. We envision this dataset as a valuable resource for advancing machine learning applications in neuro-oncology, supporting both academic research and clinical decision-support development. datasetlink: https://www.kaggle.com/datasets/briscdataset/brisc2025/
Keywords:
MRI dataset, Segmentation, Classification, Brain Tumor
1 Introduction
Brain tumors are among the most critical medical conditions, necessitating precise and timely diagnosis for effective treatment and management zhu2024sparse. Magnetic Resonance Imaging (MRI) plays a pivotal role in diagnosing and monitoring brain tumors due to its non-invasive nature and ability to provide detailed images of brain structures zou2025mmr; dayarathna2024deep; lyu2025state. Despite significant advancements in medical imaging technologies, developing automated systems for tumor detection and segmentation remains a formidable challenge zhang2025large; zhang2024challenges; rezvani2025fusionlungnet. This difficulty stems primarily from the limited availability of high-quality labeled datasets tailored for these tasks, coupled with the inherent complexity and variability in tumor appearances across patients zhang2024application; askari2025enhancing; fateh2024msdnet.
Existing brain tumor segmentation datasets, such as the BraTS menze2014multimodal, Figshare Cheng2017, and others, have significantly advanced the development of automated segmentation models. However, several limitations in these datasets drive the need for novel datasets to address emerging challenges in the field. For instance, the BraTS dataset, while comprehensive and widely used, exhibits certain constraints such as its reliance on pre-processed and standardized data that may not represent real-world variability in MRI acquisition protocols across institutions. Additionally, BraTS primarily focuses on gliomas and lacks representation of other tumor types, potentially limiting the generalizability of models trained on it ghaffari2019automated; labella2024analysis. The Figshare dataset, on the other hand, suffers from class imbalance and limited diversity in imaging conditions and patient demographics, which can restrict model robustness de2024enhanced. Moreover, annotation precision and quality remain critical issues in many publicly available datasets, where inconsistencies in labeling can adversely impact the training and evaluation of segmentation algorithmsarkar2024evolution; khajeha2025advancing; fakhima2025covsgnet; rezvani2024single. These limitations underscore the necessity of introducing a new dataset that offers balanced class distributions, multi-institutional diversity, and high-quality expert annotations to enhance the reliability and generalizability of automated brain tumor segmentation models.
To address these challenges, we present a meticulously curated brain tumor MRI dataset designed to advance research in tumor detection and segmentation. Our dataset comprises 5,000 high-resolution T1-weighted MRI images for training and 1,000 for testing, carefully selected for their suitability in visualizing brain tissue and tumor regions. Each image is accompanied by precise segmentation masks, created using advanced annotation tools and validated by radiologist and physician to ensure accuracy. The dataset focuses on three of the most common brain tumor types—Glioma, Meningioma, and Pituitary tumors—as well as a ”No Tumor” class, which is essential for developing models capable of distinguishing between healthy and tumorous brain scans. By excluding T2-weighted images and other less relevant imaging modalities, we ensure consistent data quality and minimize the risk of misclassification.
An important feature of this dataset is the inclusion of multiple imaging perspectives for each tumor, covering the Coronal, Sagittal, and Axial planes. This comprehensive approach captures a diverse range of tumor characteristics, enabling researchers to train models that generalize well across different views. Furthermore, the dataset follows a well-structured division into training and test sets, with 5000 images allocated for training and 1000 for testing, ensuring a balanced distribution of tumor types across both sets.
The potential applications of this dataset extend beyond segmentation and classification tasks. It provides a valuable benchmark for evaluating the performance of deep learning models, particularly in medical imaging. Researchers can use this dataset to explore various challenges, such as tumor boundary delineation, multi-class classification, and domain adaptation in medical imaging. Additionally, the dataset holds promise for clinical applications, including the development of decision-support systems to assist radiologists and physicians in diagnosing brain tumors with greater accuracy and efficiency.
In addition to introducing the dataset, we present a novel transformer-based model, Swin-HAFUNet, designed for brain tumor segmentation. This model adopts a hierarchical encoder-decoder architecture, leveraging Swin Transformer blocks to capture both local and global contextual information. It incorporates two key innovations: the Hierarchical Attention Fusion (HAF) module and the Contextual Bottleneck Enhancer (CBE). These components enhance the model’s ability to aggregate multi-scale features and refine semantic representations, resulting in improved segmentation performance across diverse tumor types.
The main contributions of this work are summarized as follows:
  • We introduce a large-scale, high-quality brain tumor MRI dataset with expert-annotated segmentation masks and classification labels, covering three major tumor types (glioma, meningioma, pituitary) and a no-tumor category.
  • The dataset includes multi-planar views (axial, sagittal, and coronal), enabling the development of models with robust cross-view generalization capabilities.
  • We propose Swin-HAFUNet, a lightweight yet powerful segmentation model that integrates Swin Transformer blocks, the Hierarchical Attention Fusion (HAF) module, and the Contextual Bottleneck Enhancer (CBE) for effective multi-scale feature integration.
  • Extensive experiments demonstrate that our model outperforms existing state-of-the-art methods in brain tumor segmentation, achieving the highest weighted mean IoU on the benchmark.
  • We establish strong baseline results on the proposed dataset to support future research in brain tumor analysis and medical image segmentation.
The remainder of this paper is organized as follows: Section 2 reviews related work and existing datasets in brain tumor research. Section 3 provides essential background and key medical concepts related to brain tumors and MRI imaging. Section 4 presents a detailed description of the proposed dataset, including its structure, annotation process, and imaging modalities. Section 5 offers visual demonstrations and qualitative insights. Section 6 describes the proposed Swin-HAFUNet segmentation model. Section 7 provides experimental evaluations and comparative results. Finally, Section 8 concludes the paper and outlines future research directions.
2 Related Work
In recent years, significant advancements in deep learning have driven the development of automated brain tumor diagnosis systems. A key factor in achieving high-performing models is the availability of high-quality annotated datasets, which serve as a foundation for both segmentation and classification tasks. Numerous brain tumor datasets have been introduced, each varying in imaging modalities, annotation quality, tumor types, and overall dataset size. This section provides a comprehensive review of existing brain tumor datasets, categorizing them based on their intended tasks—segmentation or classification—and highlighting their key attributes, strengths, and limitations. By examining these datasets, we aim to underscore the challenges they present and justify the need for introducing our novel dataset with refined annotations and enhanced diagnostic utility.
2.1 segmentation
When addressing brain MRI segmentation, the Brain Tumor Segmentation (BraTS) challenge remains a pivotal benchmark in the field menze2014multimodal. BraTS, which undergoes annual updates, has evolved significantly, with its latest iteration—BraTS 2024—incorporating a diverse dataset encompassing various tumor types, including gliomas, pediatric brain tumors, and brain metastases de20242024. While BraTS provides high-quality multimodal MRI data and expert-annotated tumor sub-regions, it predominantly emphasizes cases with multiple MRI sequences, such as T1, T2, T1-contrast, and FLAIR. This focus may not fully reflect real-world clinical scenarios where single-sequence scans are more commonly encountered. Additionally, BraTS’ reliance on synthetic modalities for specific cases introduces variability that might not always correlate with actual clinical imaging conditions. These limitations underscore the need for complementary datasets that encompass a broader range of imaging scenarios and real-world variability.
Another notable dataset in brain tumor segmentation is the Medical Segmentation Decathlon (MSD) antonelli2022medical. Unlike the original BraTS challenge, which primarily targets gliomas, MSD encompasses a broader spectrum of medical imaging tasks, including brain tumor segmentation. The brain MRI data in MSD originates from multiple medical centers, enhancing the dataset’s heterogeneity and, consequently, the potential generalizability of models trained on it. However, this multi-center nature introduces variability in imaging protocols, scanner types, and acquisition parameters, complicating model training and evaluation.
In addition to BraTS and MSD, the Federated Tumor Segmentation (FeTS) dataset pati2021federated represents another key resource in brain tumor segmentation. FeTS builds upon the BraTS dataset while incorporating additional clinical data from multiple healthcare institutions. It primarily focuses on gliomas, with multimodal MRI scans meticulously annotated by expert radiologists and physicians. The inclusion of multi-center data enhances the generalizability of models developed using FeTS. However, FeTS is predominantly centered on glioma cases, limiting its applicability to other tumor types, such as meningiomas and pituitary tumors. This narrow scope highlights the need for datasets that comprehensively cover a wider variety of brain tumor types for more generalized segmentation tasks.
Beyond BraTS, MSD, and FeTS, several other datasets contribute to advancing brain MRI segmentation. BrainMetShare BrainMetShare is one such dataset, comprising 156 whole-brain MRI studies with high-resolution, multi-modal pre- and post-contrast sequences from patients presenting with at least one brain metastasis. Ground-truth segmentations provided by expert radiologists and physicians accompany each study, making it a valuable resource for developing models tailored to brain metastases. Unlike glioma-focused datasets, BrainMetShare’s emphasis on metastatic brain tumors offers a complementary perspective, addressing an important clinical need.
The Open Access Series of Imaging Studies (OASIS) project OASIS also provides a significant resource for brain MRI research. While not explicitly designed for tumor segmentation, OASIS offers a large-scale collection of multimodal brain MRI scans, including longitudinal data from healthy individuals and patients with various neurological disorders. OASIS-3, in particular, features structural MRI scans acquired over time, enabling the study of disease progression. Although primarily intended for neuroscience research, OASIS’ comprehensive data can be leveraged for developing segmentation models that generalize across different brain conditions, including tumors.
Another valuable initiative is fastMRI, spearheaded by Facebook AI Research and NYU Langone Health fastMRI. While the primary objective of fastMRI is to promote advancements in MRI reconstruction, the dataset includes a substantial number of fully sampled brain MRI scans acquired on 1.5T and 3T scanners. These high-resolution scans, encompassing T1-weighted, T2-weighted, and FLAIR sequences, can be repurposed for segmentation tasks. The large scale and high diversity of the fastMRI dataset make it a valuable resource for training robust models, particularly in scenarios where high-quality input data is critical.
The LGG Segmentation Dataset LGG_MRI_Segmentation, available on Kaggle, specifically targets lower-grade gliomas by providing brain MRI scans alongside manual FLAIR abnormality segmentation masks. This dataset addresses a gap by offering focused data for a specific tumor grade, enabling the development of models tailored to low-grade glioma segmentation. Although limited in scope, its accessibility and detailed annotations make it a popular choice for researchers developing segmentation algorithms for lower-grade gliomas.
2.2 classifcation
The development of robust brain tumor classification models heavily relies on high-quality datasets that provide diverse and well-annotated medical imaging data. Several publicly available datasets have been widely used in recent research, enabling the training and evaluation of deep learning models for brain tumor classification. These datasets vary in size, imaging modalities, and tumor types, offering researchers a range of options for developing and benchmarking their models.
One of the most widely used datasets is the Figshare Cheng2017, which contains 3,064 T1-weighted contrast-enhanced MRI (CE-MRI) images categorized into three tumor types: glioma, meningioma, and pituitary tumors. This dataset has been instrumental in advancing brain tumor classification research. It provides a balanced distribution of tumor types, making it suitable for multi-class classification tasks. The dataset’s accessibility and comprehensive annotations have made it a benchmark for evaluating the performance of deep learning models, as demonstrated in studies such as Islam et al. islam2024brainnet and Balamurugan et al. ag2024robust.
Another notable dataset is one of the Kaggle Brain Tumor MRI datasets, introduced by Nickparvar Nickparvar_kaggle_brain_tumor, which includes 7,023 brain MRI images categorized into four classes: glioma, meningioma, pituitary tumors, and no tumor. This dataset is particularly valuable for its diversity and inclusion of non-tumor cases, allowing researchers to develop models capable of distinguishing between healthy and pathological brain scans. The dataset has been used in studies such as Chen et al. chen2024robust and Alanazi et al. alanazi2022brain, where it facilitated the development of transfer learning and feature fusion approaches for brain tumor classification.
The BraTS (Brain Tumor Segmentation) dataset menze2014multimodal is another critical resource, primarily focused on tumor segmentation but also widely used for classification tasks. The dataset includes multi-modal MRI scans (T1, T1c, T2, and FLAIR) with annotated tumor regions, making it suitable for both segmentation and classification tasks. The BraTS dataset has been used in studies such as Ghosal et al. ghosal2019brain, where researchers developed a Squeeze and Excitation ResNet model for brain tumor classification, achieving an accuracy of 93.83%. The dataset’s multi-modal nature allows for the exploration of complementary information from different imaging modalities, enhancing model performance.
The ”Brain Tumor” dataset gaillard2024brain is another valuable resource, providing a collection of brain MRI images with annotations for various tumor types. The dataset’s inclusion of rare tumor types and diverse imaging protocols makes it a valuable resource for developing models that can generalize across different clinical settings.
In summary, the availability of diverse and well-annotated datasets has been instrumental in advancing brain tumor classification research. These datasets continue to play a critical role in driving innovation in the field, providing the foundation for future research and clinical applications. However, despite the progress made, there is still a pressing need for new datasets that address specific challenges, such as the inclusion of rare tumor types, multi-modal imaging data, and more diverse patient demographics. Additionally, existing datasets often focus on either segmentation or classification tasks, limiting their utility for models that require simultaneous learning of both tasks. The introduction of a new dataset that supports both segmentation and classification tasks can bridge this gap, providing a more comprehensive resource for training and evaluating models. Such datasets can also incorporate advanced annotations, such as tumor sub-regions and molecular markers, which are critical for developing models that align with clinical needs. These datasets continue to play a critical role in driving innovation in the field, providing the foundation for future research and clinical applications.
3 Background and Key Concepts
3.1 Overview of Brain Tumors
Brain tumours are amongst the most fatal of all cancers aldape2019challenges. Amongst paediatric solid tumours, brain tumours are most fatal and commonly occurring aldape2019challenges. There is diversity in the types of brain tumours, including but not limited to gliomas which account for 45% of brain tumours with pituitary tumours and meningiomas accounting for 15% each senan2022early. The gold standard imaging for diagnosis of a brain tumours or brain metastases is an MRI scan with gadolinium lapointe2018primary. When possible, management is initially via surgery to remove the lesion which is then sent for histological and molecular genotype identification lapointe2018primary. Pre, intra and post-operative MRIs can be used to guide surgical resection and management lapointe2018primary; sunaert2006presurgical; fateh2022providing. Intra-operatively, functional MRIs visualise cerebrovascular activity which can be correlated with neuronal activity, aiding the surgical team. Other techniques such as cord simulation can also be used sunaert2006presurgical. In many scenarios, regardless of skill, neurosurgical reach has to be limited for safety due to the presence of many functionally important regions within the organ aldape2019challenges; lapointe2018primary. Further management includes the use of medical interventions for symptomatic management, radiotherapy and chemotherapy aldape2019challenges. The blood brain barrier poses challenges to medical interventions and chemotherapy, this barrier filters material entering the brain via circulation, limiting medical access to the brain aldape2019challenges. Localisation of the lesion and adjacent structures via MRIs can guide both surgical and radiotherapeutic planning mehndiratta2011brain. In summary, MRI scans are used in the initial diagnosis and management planning of brain tumours, including pre-surgical use and as guidance for radiotherapy mehndiratta2011brain.
3.1.1 Glioma
Gliomas are primary brain tumours, they are the most common of malignant primary brain tumours in adults lapointe2018primary. They arise from glial cells or stem cells which develop glial properties during neoplastic changes perry2016histologic. Glial cells designate a group of different cells which provide support for neurons, for example by the formation of axonal myelin sheaths jessen2004glial. For adults, the most aggressive form of gliomas, the glioblastoma, has a two year prognosis aldape2019challenges. Gliomas can be typed against the WHO 2016 classification of CNS tumours perry2016histologic.
Diffuse forms of gliomas can grow in irregular shapes, extensively infiltrating brain parenchyma perry2016histologic, this makes neurosurgical management difficult as safe maintenance of functional brain tissue is required during resection aldape2019challenges; lapointe2018primary.
3.1.2 Meningioma
Meningiomas in adults are the most common, being 30% of central nervous system tumours, whilst they are rare amongst children fathi2013meningioma. They arise from cells on the outer layer of the arachnoid mater, a part of the meninges fathi2013meningioma. The meninges is a layer in the central nervous system which encompasses the brain, cerebrospinal fluid and spinal cord dasgupta2019developmental. Though, a meningioma could arise anywhere on the meninges, 98% of meningiomas are intracranial fathi2013meningioma. Usually benign lesions, they can be slow growing fathi2013meningioma. MRI scans in conjunction with CT scan can be sued for diagnosis and treatment planning fathi2013meningioma. Treatment generally consists of neurosurgical treatment, occasionally in adjunct with radiotherapy fathi2013meningioma. Benign meningiomas can grow to a large size with pressures on the brain causing symptoms marosi2008meningioma. Non-benign meningiomas are associated with irregular shapes and tumours heterogeneity and therefore, benign meningiomas tend to be associated with regular shapes and homogeneity yan2017potential.
3.1.3 Pituitary Tumors
These are tumours originating in the pituitary gland, a small structure at the base of the brain, above the sphenoid bone bancalari2012pituitary. The pituitary gland has an essential role in growth, metabolism and reproduction asa2002pathogenesis. Due to this, pituitary tumours can cause a wide range of symptoms including but not limited to; mood disorders, diabetes mellitus, obesity, infertility and visual disturbances asa2002pathogenesis. However, only one third of these tumours are symptomatic asa2002pathogenesis. The majority of pituitary tumours are benign and when treated, treatment generally includes neurosurgical resection and radiotherapy raverot2021aggressive.
3.1.4 Non-Tumorous Conditions
There are space occupying lesions which are non tumorous in nature, they can be inflammatory, like infections, or arise from vascular abnormalities. Some examples of such space occupying lesions include abscesses, cysts, haematomas and aneurysms kelly2008neurological. Due to its high soft tissue sensitivity, MRIs can be used to aid characterisation of space occupying lesions, though surgical biopsy may still be required to finalise results reangclinico.
3.2 Magnetic resonance imaging in Brain Imaging˝
3.2.1 Anatomical planes
The main anatomical planes are coronal, transverse and sagittal planes kenhub_anatomical_terminology. These may be more simply described as a ‘front to back’ vertical plane, a ‘top to bottom’ horizontal plane, a longitudinal ‘side to side’ plane kenhub_anatomical_terminology. For a radiologist reporting an MRI scan, the above planes are available for viewing westbrook2018mri.
3.3 MRI Basics
3.3.1 MRI scans
Magnetic resonance imaging uses non hazardous electromagnetic radiation to provide images of the body’s internal structure katti2011magnetic. A computer constructs images from the information gathered by the MR scanner katti2011magnetic. These images can be viewed in sequential 2D ‘slices’ and can also be used to develop a 3D image katti2011magnetic. There are two main types of images usually provided, T1 weighted and T2 weighted images katti2011magnetic.
3.3.2 T1-Weighted Imaging
This form of imaging fat is displayed brightly, therefore it can provide detailed anatomical images of soft tissues especially in the brain katti2011magnetic.
3.3.3 T2-Weighted Imaging
In this format, the image represents water, such as CSF, brightly katti2011magnetic, due to this they can be used to detect inflammation katti2011magnetic and aid distinguish tumours from other space occupying lesions katti2011magnetic. Most of the diagnostics for space occupying lesions can be done using these two imaging modalities faehndrich2011neuroradiological.
3.3.4 Further modalities
Further MRI modalities are used by radiologists in conjunction with T1 and T2-weighted imaging to aid space occupying lesion diagnostics, these include perfusion-weighted imaging, diffusion-weighted imaging, MR spectroscopy faehndrich2011neuroradiological and FLAIR martucci2023magnetic. A radiologist generally has access to the full range of slices and may have further imaging modalities to aid diagnostic reasoning.
3.4 Challenges in Brain Tumor Diagnosis
3.4.1 Misdiagnosis Risks
As mentioned above, some tumours do not cause symptoms until reaching a certain growth marosi2008meningioma, this may lead to late clinical suspicion to warrant imaging. When available and possible, the best imaging modality are MRI scans lapointe2018primary. Even the use of neurological imaging can lead to misdiagnosis as neoplastic and non neoplastic conditions can mimic each other. As mentioned above, there are non tumorous space occupying lesions, these can be benign, meaning surgical resection and biopsy exposes many patients unnecessarily to the risks of surgery omuro2006pitfalls. There are also neoplastic brain lesions which do not appear as a space occupying lesion omuro2006pitfalls. Not neglecting T1 precontrast imaging can aid avoidance of misdiagnosis omuro2006pitfalls. Further MR imaging modalities and a thorough clinical assessment alongside some further investigations can aid in reducing errors omuro2006pitfalls. There are certain tumours which are difficult to visualised on MRI, though MRIs provide a high level of diagnostic accuracy for most tumours yan2016accuracy.
4 Dataset Description
4.1 Dataset Overview
[…]
4.2 Purpose and Objectives
The primary goal of this dataset is to overcome limitations observed in existing brain tumor datasets, such as class imbalance, lack of diversity, and annotation inconsistencies. While datasets like BraTS have driven significant advancements in glioma segmentation, their exclusive focus on gliomas and reliance on pre-processed data limit their generalizability to other tumor types and real-world scenarios. Our dataset expands the scope by incorporating multiple tumor types and includes a ”No Tumor” class to aid in broader diagnostic tasks. This addition makes the dataset highly versatile, enabling its use in applications ranging from multi-class tumor classification to binary tumor detection. The key objectives include: 1. Supporting segmentation tasks by providing accurate tumor masks. 2. Facilitating multi-class classification through balanced representation of tumor types. Axial Coronal Sagittal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 6: Samples of No Tumor across different imaging planes Original Image Initial Annotation Corrected Annotation Axial [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Coronal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Sagittal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 7: Samples of whole-region misannotations. The red area indicates regions that were initially marked as tumors but were identified by radiologist and physician as non-tumorous. Original Image Initial Annotation Regions Excluded by Radiologist and Physician (Non- Tumor) Combined Annotations Final Annotations Axial [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Coronal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Sagittal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 8: Samples of partial-region overannotations. The red area indicates regions that were initially marked as tumorous but were later identified by the radiologist and physician as non-tumorous. Original Image Initial Annotation Additional Regions Annotated by Radiologist and Physician Combined Annotations Final Annotations Axial [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Coronal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Sagittal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 9: Samples of partial-region underannotations. The purple area indicates regions that were initially marked as non-tumorous but were later identified by the radiologist and physician as tumorous.
4.3 Dataset Composition and Planar Distributions
The dataset comprises 6,000 MRI images, divided into training and testing sets, as detailed in Table 10. This structured division ensures robust evaluation metrics while providing ample data for training advanced machine learning models. For the training dataset, the total number of images across the planes is 5,000, and for testing, the total is 1,000. In addition to the class-based distribution, we provide another form of distribution—dataset composition by MRI planes. This breakdown categorizes images into Coronal, Sagittal, and Axial planes, helping to analyze how different orientations are represented in the dataset. As shown in Table 11, the distribution of different MRI planes is nearly uniform, similar to the distribution of different classes in Table 10. This balanced distribution ensures that no particular class or plane is overrepresented, which is crucial for preventing model bias and improving generalization.
4.4 Data Source and Preprocessing
This dataset was derived from ”Brain Tumor MRI Dataset” nickparvar2021brain that combined data from three prominent sources: Figshare Cheng2017, SARTAJ bhuvaji2021brain, and Br35H hamada2020br35h. The original dataset included 7,023 MRI images across four classes: Glioma, Meningioma, No Tumor, and Pituitary. The ”No Tumor” class images were specifically sourced from the Br35H dataset. During preprocessing, several steps were undertaken to ensure the quality and consistency of our final dataset: Separation of T1 Images: Only T1-weighted MRI images were retained to maintain uniformity in imaging modality. Quality Control: Images with incorrect or inconsistent labels were identified and removed with the assistance of radiologist and physician. Standardization: Images were resized and margins were adjusted to improve model accuracy during training.
4.5 Imaging Details
All images in the dataset are T1-weighted contrast-enhanced MRI scans, selected specifically from the ”Brain Tumor MRI Dataset” (Kaggle) nickparvar2021brain. Although the original dataset included some T2-weighted images, we exclusively selected T1-weighted scans for their superior ability to highlight tumor boundaries effectively. Another notable characteristic of this dataset is the length of MRI sequences. While typical brain MRI studies often consist of longer sequences, the majority of sequences in this dataset were notably short, ranging from 1 to 5 images per sequence. Sequences with only one image were excluded, as even experienced radiologists and physicians found it challenging to identify tumors accurately in these cases.
4.6 Annotation Process
The dataset underwent a meticulous annotation and review process to ensure accuracy and reliability. Annotation was performed using the AnyLabeling tool nguyen2024anylabeling, which facilitated efficient and precise delineation of tumor regions. Each image was reviewed and edited multiple times with the input of a certified physician and radiologist. Key steps in the annotation process included: Tumor Mask Refinement: Using AnyLabeling nguyen2024anylabeling, points of interest for tumors were iteratively refined to ensure precise segmentation masks. Class Correction: Images incorrectly classified as ”No Tumor” in the original dataset were re-evaluated and removed from this class if necessary. Consensus Reviews: Discrepancies in annotation were resolved collaboratively by the physician and radiologist.
5 Visual Demonstrations of the Dataset
This section provides an in-depth exploration of the dataset through visual examples and analytical discussions, illustrating its structure, composition, and inherent challenges. These demonstrations aim to deepen understanding of the dataset’s unique features while highlighting its potential applications in advanced segmentation and classification tasks.
5.1 Overview of Classes
The dataset encompasses four distinct classes: ”Glioma”, ”Meningioma”, ”Pituitary” tumors, and ”No Tumor”. As detailed in Section 3.1, each class presents unique characteristics and complexities. This subsection offers representative visual examples from each class, including raw MRI scans alongside their annotated tumor masks, emphasizing the diversity and precision of the dataset. Glioma: Gliomas are irregularly shaped and often infiltrate surrounding tissues, presenting significant challenges for precise boundary definition. These complexities require robust segmentation techniques to capture their variable morphology. As shown in Table 12, gliomas exhibit irregular and diffuse growth patterns, which are highlighted through annotated tumor masks. Meningioma: Meningiomas are generally well-circumscribed and homogeneous, making them easier to segment. However, their proximity to sensitive regions such as the meninges can complicate diagnostic tasks. An example of a meningioma and its segmentation mask is presented in Table 13, illustrating the clarity of its boundaries. Pituitary Tumors: Located at the base of the brain near critical structures like the optic chiasm, pituitary tumors demand careful delineation to avoid diagnostic errors. As shown in Table 14, the segmentation accurately captures the tumor’s boundaries without encroaching on adjacent critical regions. No Tumor: This control class is integral for training models to distinguish healthy scans from those with abnormalities. Including ”No Tumor” cases enhances the dataset’s robustness for binary and multi-class classification. Table 15 illustrates an example of a healthy brain scan with no abnormalities.
5.2 Tumor Mask and Annotation Quality
Achieving accurate tumor segmentation required a meticulous process of iterative reviews and refinements, conducted in close collaboration with a physician and a radiologist. This collaborative effort was crucial in ensuring that the final annotations accurately reflected the true tumor boundaries, minimizing errors and improving the overall quality of the dataset. In some cases, regions initially annotated as tumors were later identified by physician and radiologist as non-tumorous. These corrections were essential to avoid false positives that could mislead model training. An example of such a case is shown in Table 16, where an area initially believed to be a tumor was excluded from the final annotation after expert review. In other instances, certain areas that were mistakenly included as part of the tumor region were refined based on radiologist and physician feedback. These areas, though visually similar to tumor tissue, were determined to be non-tumorous upon closer examination. As illustrated in Table 17, the removal of these incorrect segments resulted in more precise tumor masks and enhanced the reliability of the dataset. Conversely, there were cases where genuine tumor regions had been overlooked during the initial annotation process. With input from the Physician and radiologist, these missing regions were added to the annotations, ensuring that the masks comprehensively captured all tumor areas. Table 18 demonstrates an example of such an adjustment, where previously unannotated tumor segments were correctly incorporated into the final mask.
5.3 Challenges of No Tumor Conditions
Non-tumorous conditions frequently mimic tumors in MRI scans, posing significant challenges not only for classification tasks but also for segmentation. This section visualizes examples of these conditions and compares them with actual tumors to highlight their distinctions. Brain lesions, for instance, often resemble tumors both in shape and intensity. These similarities can lead to misclassification as well as erroneous segmentation of the lesion as a tumor. Cysts are another condition that can complicate both segmentation and classification. Typically fluid-filled and round, cysts may be mistaken for tumors during segmentation tasks due to their well-defined boundaries. Calcifications, which appear as bright regions on MRI scans, can similarly lead to errors in both classification and segmentation. While their growth patterns differ from tumors, their intensity can cause segmentation models to incorrectly label them as tumorous regions. These challenges underscore the critical importance of radiologist and physician expertise in ensuring accurate segmentation and classification of such conditions. The examples provided highlight the need for robust models capable of distinguishing these non-tumorous conditions from actual tumors in both classification and segmentation tasks. Refer to caption Figure 1: the overview of the Swin-HAFUNet
6 Proposed method
6.1 Overview
Although the original dataset was designed primarily for classification tasks, through close collaboration with physicians and radiologists, we have extended it by providing high-quality expert annotations delineating tumor regions, thereby creating a new segmentation benchmark. This enriched dataset motivated the development of an effective segmentation model tailored to accurately localize and delineate brain tumors in MRI scans. In this paper, we present a novel transformer-based architecture for accurate and efficient tumor segmentation in brain MRI scans. The overall framework adopts an encoder-decoder structure, where the encoder extracts multi-scale semantic representations from the input image, and the decoder progressively reconstructs the segmentation map using enhanced contextual features. The encoder is built upon a hierarchical Swin Transformer backbone, which efficiently captures both local and global dependencies through shifted window-based self-attention. To further refine the extracted features, we use the Contextual Bottleneck Enhancer (CBE), which enriches feature representations using a sequence of lightweight yet effective operations, including shifted MLPs and gated encoding units. To preserve high-resolution semantic details during decoding, we design a lightweight decoder that includes Adaptive Context Aggregator blocks, which adaptively fuse local and global context from the encoder outputs. Additionally, we propose a Hierarchical Attention Fusion (HAF) module that integrates multi-scale features from different encoder levels through a combination of Swin Transformer blocks and deformable convolutions, allowing the model to capture hierarchical dependencies and spatial variations effectively. Refer to caption Figure 2: Hierarchical Attention Fusion module Together, these components enable our model to achieve robust segmentation performance while maintaining computational efficiency. The complete architecture is illustrated in Figure 4.
6.2 Encoder Architecture
The encoder of the proposed segmentation model is designed to extract rich hierarchical features from brain MRI scans using a multi-stage Swin Transformer-based backbone. It begins with a Patch Partition module, which splits the input image into non-overlapping patches. These patches are then flattened and mapped to a fixed-dimensional embedding space through a Linear Embedding layer. Following this, the encoder comprises three repeated stages, each consisting of two Swin Transformer Blocks followed by a Patch Merging layer. The Swin Transformer blocks utilize a window-based self-attention mechanism with a shifted window strategy, allowing the model to effectively capture local and non-local dependencies with reduced computational cost. The patch merging operation downsamples the spatial resolution while increasing the feature dimensionality, forming a hierarchical representation. This hierarchical structure enables the encoder to progressively capture multi-scale semantic information, which is crucial for segmenting tumors of varying sizes and shapes. The feature maps from different stages are later passed to the HAF modules, enabling multi-level feature interaction and refinement in the decoding process.
6.3 Hierarchical Attention Fusion (HAF) Module
The HAF module is designed to effectively integrate encoder and decoder features at each resolution level, enhancing the model’s capacity to capture both low-level spatial detail and high-level semantic context. At each stage of the decoder, the HAF module takes two inputs: $x_{\text{skip}}$: The skip connection feature map from the encoder. $x_{\text{decoder}}$: the upsampled feature map from the previous decoder layer. The decoder feature $x_{\text{decoder}}$ is first passed through a Swin Transformer Block to refine contextual dependencies and enhance representation. As shown in Equation 4, the refined feature is then concatenated with the corresponding encoder feature $x_{\text{skip}}$ along the channel dimension. $x_{\text{cat}}=\text{Concat}(x_{\text{skip}},\ \text{Swin}(x_{\text{decoder}}))$ (1) This concatenated feature map $x_{\text{cat}}$ is then projected through a $1\times 1$ convolutional layer (Equation 5). $x_{\text{out}}=\text{Conv}_{1\times 1}(x_{\text{cat}})$ (2) The resulting output $x_{\text{out}}$ maintains the same spatial resolution as $x_{\text{decoder}}$ and serves as the input for the next step in the decoder. This fusion strategy allows the network to maintain fine spatial information while enriching the semantic features through transformer-based attention. The Architecture of HAF is shown in Figure 5.
6.4 Contextual Bottleneck Enhancer (CBE)
The CBE employs a Tokenized MLP Block to enhance feature representations efficiently at the bottleneck stage. This module captures long-range dependencies while maintaining computational efficiency. Refer to caption Figure 3: Structure of the Adaptive Context Aggregator module. The process begins with a spatial shift along the width axis, followed by a linear projection to produce token embeddings. These tokens pass through a shifted MLP layer (across width), then through a depth-wise convolution and GELU activation, which introduces non-linearity and encodes positional information. The output is then processed by another shifted MLP along the height axis. A residual connection adds the original input tokens to the output of the second MLP, and the result is normalized using Layer Normalization (LN). This sequence can be summarized in Equation 6. 1. $X_{\text{shift}}^{W}$ $=\text{Shift}_{W}(X)$ (3) 2. $T_{W}$ $=\text{MLP}_{W}(X_{\text{shift}}^{W})$ 3. $Y$ $=\text{GELU}(\text{DWConv}(T_{W}))$ 4. $Y_{\text{shift}}^{H}$ $=\text{Shift}_{H}(Y)$ 5. $T_{H}$ $=\text{MLP}_{H}(Y_{\text{shift}}^{H})$ 6. $Z$ $=\text{LN}(T_{H}+T)$ where $T$ represents the original tokenized input. This design enables the block to model spatial context in both directions and fuse the information effectively via residual learning.